# Identifying and Evaluating Inactive Heads in Pretrained LLMs

### v2

@@ 2 Background and Related Work
The multi-head self-attention mechanism (Vaswani, 2017) allows multiple attention heads to operate in parallel, each focusing on different aspects of the input sequence. For an input of NN tokens with dimensionality $d_{m}$, learned linear projections transform queries, keys, and values ($\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d_{m}}$) into lower-dimensional representations: $d_{k}$-dimensional queries/keys and $d_{v}$-dimensional values, specific to each of the $h$ heads. The self-attention operation within each head computes attention weights, defined as $\mathbf{A}=\mathrm{softmax}(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}})$, where $\mathbf{A}\in[0,1]^{N\times N}$ and each row sums to 1, which are then applied to the value vectors. Another way to view the output of an attention head is that, for every position in the sequence, we compute a convex combination of value vectors, weighted by the corresponding row of $\mathbf{A}$. The outputs of all heads are concatenated and passed through a final linear projection to produce the module’s output. The attention weights, $\mathbf{A}$, are one of the few features we can visualize to get a sense of how pretrained LLMs work. For a token sequence, high attention weights (close to 11) reveal which tokens are considered most relevant. Thus, it is surprising that attention can concentrate on initial tokens, called ”attention sinks”, as shown in App. Figure 10.
@@
Guo et al. (2024a) propose the idea that attention heads can be either active or dormant. They perform a model intervention study on 33 attention heads of Llama-2-7B Base, where a specific attention head output is zeroed out, and the difference in loss as a result of the intervention is measured. They find that the difference in loss can depend on whether input text is from Wikipedia or GitHub; a head that is dormant on Wikipedia samples does not change the loss when the head output is zeroed. In our work, we go beyond analyzing one attention head at a time and propose to zero out all “dormant heads” while evaluating models on a real-world benchmark task. If “dormant heads” are truly inactive, then zeroing them out should have little effect on model performance. We also propose new score functions that measure different definitions of inactive. Unlike Guo et al. (2024a) and Gu et al. (2025) who measure only attention weights, our score functions also consider value vectors and head outputs. By considering the range of ways a head can be inactive, we present a more accurate picture of inactive attention heads in pretrained models. A better understanding of inactive attention heads has the potential to be used for efficiency (Li et al., 2021), KV cache reduction (Liu et al., 2023), and compression (Ge et al., 2024).
@@ 4.1 Setup
We download $14$ pretrained models using Hugging Face transformers (Wolf et al., 2020): Llama-3.1-8B, Llama-3.1-8B-Instruct (Dubey et al., 2024), Llama-3.2-3B, Llama-3.2-3B-Instruct (Meta, 2024), OLMo-2-1124-7B, OLMo-2-1124-7B-SFT, OLMo-2-1124-7B-DPO, OLMo-2-1124-7B-Instruct (OLMo et al., 2024), Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-3B, Qwen2.5-7B, Qwen2.5-7B-Instruct, and Qwen2.5-14B (Yang et al., 2024). We refer to models from the same organization as belonging to the same “model family.” We evaluate models on MMLU (Hendrycks et al., 2021) (55-shot) in our model intervention experiments. For datasets that do not use an LLM-as-a-Judge, MMLU is the multiple-choice (MC) benchmark with the highest Spearman correlation with ChatBot Arena (Li et al., 2023; Chiang et al., 2024), and thus makes it the most appropriate choice for aligning with real-world chatbot preferences while keeping evaluation consistent and tractable. All evaluations are performed using lm-evaluation-harness (Gao et al., 2024). We also use FineWeb-Edu (Lozhkov et al., 2024), which is a pretraining dataset of educational webpages. We randomly truncate FineWeb-Edu training sequences to between $10$ and $3000$ tokens to exhibit a variety of sequence lengths. Additional model and dataset details are in Section A.9.
@@ 4.3 Verifying Inactive Heads through Model Interventions
In Figure 4, we measure how MMLU accuracy changes as we zero out heads selected by different score functions. For brevity, we present one model from each model family: Llama-3.1-8B-Instruct, OLMo-2-1124-7B-Instruct, and Qwen2.5-3B. To do this, we vary the threshold of each score function such that identified heads are at most $30\%$ of the model’s heads. We plot the top-3 score functions for every model in Figure 4, but complete results for all models and score functions can be found in App. Figure 6. In Table 2, we quantify the percent of heads that can be zeroed while keeping accuracy to within $1\%$ of baseline. We find that, for 8 out of 14 models, the Avg Head Output Norm (LN) score function is best at identifying the most attention heads that when zeroed, maintain accuracy to within $1\%$ of baseline. For 13 out of 14 models, AHON (LN) ranks in the top-33 scoring functions. Avg Weight of First Token underestimates the number of inactive heads.
@@ 4.4 Score Distributions for Studying Attention Behaviors
We use the Qwen2.5 models (Yang et al., 2024) to understand how model scale affects attention head behaviors. There are 55 model types ranging from 0.50.5B parameters to 1414B parameters, but while the models share the same pre-training data, the exact number of training tokens each model was exposed to is not publicly documented. Assuming model scale is the only variable, it appears that model scale impacts attention head behaviors. As Qwen2.5 models grow in size, they are quite similar to one another, until the 1414B scale. For example, on most score functions, Wasserstein distance is near-zero, indicated by a clear dark region that corresponds to inter-model similarity. But at the 14B scale, there appears to be only 2 score function distribution where Qwen2.5-14B is similar to other models in the Qwen2.5 family: Avg Value Vector Norm and its layer normalized version. This may suggest that larger models learn to specialize their heads in different ways. The similarity of score distributions between the Llama-3 models we consider may have more to do with their training objective as opposed to training data. In particular, because Llama-3.2 was created from pruning Llama-3.1-8B, then performing knowledge distillation, these models are trained to be similar (Meta, 2024).
@@ Appendix A Appendix
Figure 6: Complete Results of Section 4.3. For $14$ models, we consider $13$ scoring functions, and evaluate each at 77 distinct thresholds $\tau$.
@@ A.2 Score Distributions on MMLU
Figure 8: Wasserstein Distance of Head Score Distributions between Models. For $13$ scoring functions, we plot a matrix of Wasserstein distances between models.
@@ A.6 Redundancy of attention patterns
For every model in Figure 11, we also show the top 22 principal components across all attention matrices (of which there are $N_{\rm layer}\times N_{\rm heads}$). It is surprising to see a few principal components capturing more than half of the observed variance of attention matrices, for all models.

### v3

@@ 2 Background and Related Work
The multi-head self-attention mechanism (Vaswani, 2017) allows multiple attention heads to operate in parallel, each focusing on different aspects of the input sequence. For an input of $N$ tokens with dimensionality $d_{m}$, learned linear projections transform queries, keys, and values ($\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d_{m}}$) into lower-dimensional representations: $d_{k}$-dimensional queries/keys and $d_{v}$-dimensional values, specific to each of the $h$ heads. The self-attention operation within each head computes attention weights, defined as $\mathbf{A}=\mathrm{softmax}(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}})$, where $\mathbf{A}\in[0,1]^{N\times N}$ and each row sums to 1, which are then applied to the value vectors. Another way to view the output of an attention head is that, for every position in the sequence, we compute a convex combination of value vectors, weighted by the corresponding row of $\mathbf{A}$. The outputs of all heads are concatenated and passed through a final linear projection to produce the module’s output. The attention weights, $\mathbf{A}$, are one of the few features we can visualize to get a sense of how pretrained LLMs work. For a token sequence, high attention weights (close to $1$) reveal which tokens are considered most relevant. Thus, it is surprising that attention can concentrate on initial tokens, called ”attention sinks”, as shown in App. Figure 10.
@@
Guo et al. (2024a) propose the idea that attention heads can be either active or dormant. They perform a model intervention study on $3$ attention heads of Llama-2-7B Base, where a specific attention head output is zeroed out, and the difference in loss as a result of the intervention is measured. They find that the difference in loss can depend on whether input text is from Wikipedia or GitHub; a head that is dormant on Wikipedia samples does not change the loss when the head output is zeroed. In our work, we go beyond analyzing one attention head at a time and propose to zero out all “dormant heads” while evaluating models on a real-world benchmark task. If “dormant heads” are truly inactive, then zeroing them out should have little effect on model performance. We also propose new score functions that measure different definitions of inactive. Unlike Guo et al. (2024a) and Gu et al. (2025) who measure only attention weights, our score functions also consider value vectors and head outputs. By considering the range of ways a head can be inactive, we present a more accurate picture of inactive attention heads in pretrained models. A better understanding of inactive attention heads has the potential to be used for efficiency (Li et al., 2021), KV cache reduction (Liu et al., 2023), and compression (Ge et al., 2024).
@@ 4.1 Setup
We download $14$ pretrained models using Hugging Face transformers (Wolf et al., 2020): Llama-3.1-8B, Llama-3.1-8B-Instruct (Dubey et al., 2024), Llama-3.2-3B, Llama-3.2-3B-Instruct (Meta, 2024), OLMo-2-1124-7B, OLMo-2-1124-7B-SFT, OLMo-2-1124-7B-DPO, OLMo-2-1124-7B-Instruct (OLMo et al., 2024), Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-3B, Qwen2.5-7B, Qwen2.5-7B-Instruct, and Qwen2.5-14B (Yang et al., 2024). We refer to models from the same organization as belonging to the same “model family.” We evaluate models on MMLU (Hendrycks et al., 2021) ($5$-shot) in our model intervention experiments. For datasets that do not use an LLM-as-a-Judge, MMLU is the multiple-choice (MC) benchmark with the highest Spearman correlation with ChatBot Arena (Li et al., 2023; Chiang et al., 2024), and thus makes it the most appropriate choice for aligning with real-world chatbot preferences while keeping evaluation consistent and tractable. All evaluations are performed using lm-evaluation-harness (Gao et al., 2024). We also use FineWeb-Edu (Lozhkov et al., 2024), which is a pretraining dataset of educational webpages. We randomly truncate FineWeb-Edu training sequences to between $10$ and $3000$ tokens to exhibit a variety of sequence lengths. Additional model and dataset details are in Section A.9.
@@ 4.3 Verifying Inactive Heads through Model Interventions
In Figure 4, we measure how MMLU accuracy changes as we zero out heads selected by different score functions. For brevity, we present one model from each model family: Llama-3.1-8B-Instruct, OLMo-2-1124-7B-Instruct, and Qwen2.5-3B. To do this, we vary the threshold of each score function such that identified heads are at most $30\%$ of the model’s heads. We plot the top-3 score functions for every model in Figure 4, but complete results for all models and score functions can be found in App. Figure 6. In Table 2, we quantify the percent of heads that can be zeroed while keeping accuracy to within $1\%$ of baseline. We find that, for 8 out of 14 models, the Avg Head Output Norm (LN) score function is best at identifying the most attention heads that when zeroed, maintain accuracy to within $1\%$ of baseline. For 13 out of 14 models, AHON (LN) ranks in the top-$3$ scoring functions. Avg Weight of First Token underestimates the number of inactive heads.
@@ 4.4 Score Distributions for Studying Attention Behaviors
We use the Qwen2.5 models (Yang et al., 2024) to understand how model scale affects attention head behaviors. There are $5$ model types ranging from 0.50.5B parameters to 1414B parameters, but while the models share the same pre-training data, the exact number of training tokens each model was exposed to is not publicly documented. Assuming model scale is the only variable, it appears that model scale impacts attention head behaviors. As Qwen2.5 models grow in size, they are quite similar to one another, until the 1414B scale. For example, on most score functions, Wasserstein distance is near-zero, indicated by a clear dark region that corresponds to inter-model similarity. But at the 14B scale, there appears to be only 2 score function distribution where Qwen2.5-14B is similar to other models in the Qwen2.5 family: Avg Value Vector Norm and its layer normalized version. This may suggest that larger models learn to specialize their heads in different ways. The similarity of score distributions between the Llama-3 models we consider may have more to do with their training objective as opposed to training data. In particular, because Llama-3.2 was created from pruning Llama-3.1-8B, then performing knowledge distillation, these models are trained to be similar (Meta, 2024).
@@ Appendix A Appendix
Figure 6: Complete Results of Section 4.3. For $14$ models, we consider $13$ scoring functions, and evaluate each at 77 distinct thresholds $\tau$ for each.
@@ A.2 Score Distributions on MMLU
Figure 8: Wasserstein Distance of Head Score Distributions between Models. For each of $13$ score functions, we plot a matrix of Wasserstein distances between score distributions of different models.
@@ A.6 Redundancy of attention patterns
For every model in Figure 11, we also show the top $2$ principal components across all attention matrices (of which there are $N_{\rm layer}\times N_{\rm heads}$). It is surprising to see a few principal components capturing more than half of the observed variance of attention matrices, for all models.